OpenStudy (anonymous):

A cylindrical tank with radius 3 m is being filled with water at a rate of 4 m3/min. How fast is the height of the water increasing?

OpenStudy (anonymous):

\[ A = \pi r^2h \]Can you differentiate?

OpenStudy (anonymous):

uhm i mean i can plug the numbers in and get the answer but i don't know which parts to differentiate?

OpenStudy (anonymous):

Think about it like this: \[ A(t) = \pi \times [r(t)]^2 \times [h(t)] \]

OpenStudy (anonymous):

and then plug in the variables?

OpenStudy (anonymous):

You want to use the chain rule to differentiate.

OpenStudy (anonymous):

how do you tell when to use the chain rule?

OpenStudy (anonymous):

so would my answer be 1728pi?

OpenStudy (anonymous):

Because we are differentiating functions of functions.

OpenStudy (anonymous):

nevermind i know thats the wrong answer

OpenStudy (anonymous):

i dont know that to do

OpenStudy (anonymous):

We want to use: \[ \frac{dA}{dt} = \frac{\partial A}{\partial r}\frac{dr}{dt} + \frac{\partial A}{\partial h}\frac{dh}{dt} \]

OpenStudy (anonymous):

what is that weird symbol?

OpenStudy (anonymous):

Basically, \[ A(t) = f(r(t),h(t)) \]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Hmmm, well \(\partial \) is used for partial derivatives. Perhaps you have not learned about them yet?

OpenStudy (anonymous):

i havent learned that yet

OpenStudy (anonymous):

In that case, we'll have to use a bit more complicated method.

OpenStudy (anonymous):

Okay, so this is what we can do... \[ \frac{d}{dt}\pi r^2 h = \left(\frac{d}{dt} \pi r^2 \right)h +\pi r^2 \left(\frac{d}{dt}h\right) \]

OpenStudy (anonymous):

Then we need to simplify: \[ \frac{d}{dt} \pi r^2 = \pi 2r \frac{d}{dt}r \]

OpenStudy (anonymous):

Tell me, does this make sense at all to you so far?

OpenStudy (anonymous):

The chain rule tells us that: \[ \frac{d}{dt} \pi r^2 = \frac{d}{dr} \pi r^2 \frac{d}{dt}r \]And we know: \[ \frac{d}{dr}\pi r^2 = 2\pi r \]Which means \[ \frac{d}{dr} \pi r^2 \frac{d}{dt}r = 2\pi r \frac{d}{dt}r \]

OpenStudy (anonymous):

okay i understand that so far like it makes sense... but then which equation do we place the numbers into

OpenStudy (anonymous):

would it be the 2pir d/dt r

OpenStudy (anonymous):

Hold on. So up above we have: \[ \frac{d}{dt}\pi r^2 h = \left(\frac{d}{dt} \pi r^2 \right)h +\pi r^2 \left(\frac{d}{dt}h\right) \] And this equals: \[ = \left(2\pi r \frac{d}{dt}r \right)h +\pi r^2 \left(\frac{d}{dt}h\right) \]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Hmmm, hold on... I think I made things more complicated than necessary... It seems that \(\frac{dr}{dt}=0\) from the start....

OpenStudy (anonymous):

So actually, we just have to do: \[ A = \pi r^2h \] And we will get: \[ \frac{dA}{dt} = \pi r^2 \frac{dh}{dt} \]

OpenStudy (anonymous):

Since \(r\) is constant with respect to time, all my other stuff wasn't necessary. This problem is actually very simple. They already tell us \(r=3\text{m}\) and \(dA/dt = 4\text{m}^3/\text{min}\)

OpenStudy (anonymous):

You just solve for \(dh/dt\).

OpenStudy (anonymous):

so then would my answer be 4-9pi?

OpenStudy (anonymous):

this is so frustrating i really dont know what to do

OpenStudy (anonymous):

No, you don't want to use subtraction to solve for \(dh/dt\).

OpenStudy (anonymous):

Since it is being multiplied, you want to divide out the other factor.

OpenStudy (anonymous):

so its then 4/9pi

OpenStudy (anonymous):

thank you so much